OpenStudy (anonymous):

evaluate following piecewise defined function @ x=-3,x=1/4, and x=3 f(x)3^rt9x for x<0 8x for 01/2

OpenStudy (anonymous):

wow am I lost...

OpenStudy (campbell_st):

ok... so you have 3 functions with differing domains. I am unsure of the 1st function but you need to substitute x = -3 as x<0 next you need to find f(1/4) so you use f(x) = 8x since 0<= x <= 1/2 and lastly f(3) so you use 2/3 x since x>1/2

OpenStudy (anonymous):

???

OpenStudy (campbell_st):

so you can't find f(1/4) when f(x) = 8x?

OpenStudy (anonymous):

2

OpenStudy (campbell_st):

thats good now f(3) uses the last function f(x) = 2/3 x becuase x > 1/2 can you now find f(3)

OpenStudy (anonymous):

2

OpenStudy (campbell_st):

ok... so the conditions to the right of each function, determine which to use... now for x = -3 you need to find f(-3) which uses the 1st function, that I can't understand, because x < 0

OpenStudy (anonymous):

=-3

OpenStudy (campbell_st):

well looks like you have a solution. hope it helps

OpenStudy (anonymous):

f(x)=\[\sqrt[3]{9x}\]

OpenStudy (anonymous):

that is the 1st one. But what do I do with these numbers